### v1

# How Good is Post-Hoc Watermarking With Language Model Rephrasing?

Abstract
Generation-time text watermarking embeds statistical signals into text for traceability of AI-generated content. We explore post-hoc watermarking where an LLM rewrites existing text while applying generation-time watermarking, to protect copyrighted documents, or detect their use in training or RAG via watermark radioactivity. Unlike generation-time approaches, which is constrained by how LLMs are served, this setting offers additional degrees of freedom for both generation and detection. We investigate how allocating compute (through larger rephrasing models, beam search, multi-candidate generation, or entropy filtering at detection) affects the quality-detectability trade-off. Our strategies achieve strong detectability and semantic fidelity on open-ended text such as books. Among our findings, the simple Gumbel-max scheme surprisingly outperforms more recent alternatives under nucleus sampling, and most methods benefit significantly from beam search. However, most approaches struggle when watermarking verifiable text such as code, where we counterintuitively find that smaller models outperform larger ones. This study reveals both the potential and limitations of post-hoc watermarking, laying groundwork for practical applications and future research.
[…]
1 Introduction
Post-hoc text watermarking inserts an algorithmically-detectable signal into an existing text while remaining imperceptible to readers. It serves several goals such as copyright protection or traitor tracing. Early approaches relied on hand-crafted modifications: synonym substitutions (Topkara et al., 2006c; Shirali-Shahreza and Shirali-Shahreza, 2008), grammatical transformations (Topkara et al., 2006b, a) or morphosyntactic alterations (Meral et al., 2009). Recent methods based on deep neural networks follow a post-hoc paradigm where the input is the original text and the output is the watermarked text, using one watermark embedder and one watermark extractor (Abdelnabi and Fritz, 2021), similarly to state-of-the-art post-hoc watermarking approaches for images, audio, and video. However, these methods are not effective, suffering from low capacity or unreliable detection, because they are highly constrained or easily broken by reversing the edits or synonyms.
With the emergence of large language models (LLMs), their popularization through ChatGPT (OpenAI, 2022), and growing concerns about potential risks (Crothers et al., 2022; Weidinger et al., 2022), generation-time text watermarking algorithms have been introduced to help detect AI-generated text (Aaronson and Kirchner, 2023; Kirchenbauer et al., 2023a). These algorithms have been deployed at scale, for instance in Google’s Gemini with SynthID (Dathathri et al., 2024). Most LLM watermarks alter the next token selection, for example, by promoting a specific set of tokens depending on previous tokens and a secret key. Detection of the watermark in a text is then performed through a theoretically grounded statistical test that provides rigorous guarantees over the false positive rates. In addition, these methods achieve high watermark power while adding minimal latency, by leveraging the entropy of the LLM to embed the signal. However, in contrast with previous approaches, these methods require control at generation time and therefore cannot be applied to existing text.
A natural idea to extend these methods to post-hoc watermarking is to employ a paraphrasing LLM to re-generate the text, allowing the watermark to be injected at inference time without further model training. This approach has been explored in data protection literature (Jovanović et al., 2025; Sander et al., 2025; Rastogi et al., 2025; Zhang et al., 2025; Lau et al., 2024). For example, “watermark radioactivity” exploits the fact that watermark signals leave detectable traces when watermarked text is used by another model, enabling active training and context membership inference (Sander et al., 2025; Jovanović et al., 2025). However, to our knowledge, there is no thorough evaluation of how post-hoc watermarking through LLM paraphrasing performs across different data domains, such as prose versus verifiable text like code. Moreover, unlike generation-time watermarking where watermark embedding should not delay the generation, post-hoc watermarking allows trading additional compute for a better quality-detectability trade-off. This opens up several axes to explore: model size, watermarking method, decoding strategy (and even generating multiple candidates), as well as detection strategy.
Figure 1: Post-hoc text watermarking through watermarked LLM rephrasing. We do empirical evaluations and analyze detection power, semantic fidelity, and correctness according to different design choices such as watermark scheme and available compute (through the paraphrasing model and the decoding strategy).
To address this gap, in this paper, we provide a comprehensive evaluation of post-hoc watermarking. We find that current state-of-the-art LLM watermarking schemes applied in a post-hoc setting perform well on open-ended text, such as Wikipedia articles and books, achieving high detection power while preserving fidelity. However, these methods are less effective on verifiable text like code, where the requirement to preserve correctness severely constrains paraphrasing freedom. Regarding design choices, we find that larger models better preserve semantics, while smaller models are necessary to effectively hide strong watermarks. Furthermore, the simplest Gumbel-max approach (Aaronson and Kirchner, 2023) dominates the Pareto frontier when classical random sampling is used. For other watermarking methods, we find that beam search decoding significantly improves the quality-detectability trade-off.
In short, our main contributions are:
  • A comprehensive evaluation: We conduct the first large-scale study of post-hoc watermarking through LLM rephrasing (Figure 1) across diverse domains, demonstrating that while current methods are effective for open-ended text, they struggle with verifiable formats like code.
  • An analysis of design strategies: We isolate the impact of model size, decoding, and watermarking schemes. For instance, we identify Gumbel-max as the robust Pareto-optimal choice (Figure 2).
  • An open-source research framework: We release an easily modifiable codebase to facilitate research for post-hoc watermarking techniques.
2 Related work
2.1 Post-Hoc Text Watermarking
Early text watermarking altered text characteristics like characters or spacing (Brassil et al., 1995). Other methods modify grammatical or syntactical structures via pre-established rules (Topkara et al., 2005), including synonym substitution (Topkara et al., 2006c) and word reordering through passivization or topicalization (Topkara et al., 2006b, a; Meral et al., 2009). Text steganography follows similar principles (Winstein, 1998; Chapman et al., 2001; Bolshakov, 2004; Shirali-Shahreza and Shirali-Shahreza, 2008; Chang and Clark, 2014; Xiang et al., 2017). These edit-based systems typically exhibit low robustness and payload, e.g., 1-2 bits per sentence (Wilson and Ker, 2016). Deep learning methods have since been used for this task, for instance, with masked language models for steganography (Ueoka et al., 2021), infilling models (Yoo et al., 2023a), neural lexical substitution (Qiang et al., 2023), or encoder-decoder architectures (Abdelnabi and Fritz, 2021; Zhang et al., 2024; Xu et al., 2024).
2.2 Generation-Time Watermarking With Large Language Models
The first watermarks for machine-generated text date back to a method presumably used in Google Translate to filter automated translations from future training data (Venugopal et al., 2011). For LLM-generated text, two concurrent approaches appeared shortly after the release of ChatGPT. Kirchenbauer et al. (2023a) bias a subset of the vocabulary, while Aaronson and Kirchner (2023) alter the sampling via the Gumbel trick. Both use pseudorandom seeds generated from a secret key and preceding tokens, enabling lightweight detection through statistical tests without access to the model.
Extensions include improved tests and multi-bit watermarking (Fernandez et al., 2023; Yoo et al., 2023b, 2024; Qu et al., 2024), position-dependent seeds (Christ et al., 2023; Kuditipudi et al., 2023), low-entropy optimizations (Lee et al., 2023; Christ et al., 2023; Huang et al., 2023), and semantic watermarks for better robustness (Liu et al., 2023; Liu and Bu, 2024; Fu et al., 2024; Hou et al., 2023, 2024). DiPMark (Wu et al., 2023) provides distortion-free Green-Red watermarks, and MorphMark (Wang et al., 2025) adaptively adjusts watermark strength based on the green token probability mass. WaterMax (Giboulot and Furon, 2024) generates several chunks of tokens from the original LLM distribution and selects the outputs with high watermark scores, which ensures that the LLM distribution is preserved. SynthID-Text (Dathathri et al., 2024) deploys tournament-based sampling in Google Gemini. Toolkits have also been introduced to benchmark these methods (Piet et al., 2023; Pan et al., 2024).
We provide a comprehensive description of the schemes evaluated in this work in subsection 6.2.
2.3 Post-Hoc LLM Watermarks for Data Protection
Recent works apply LLM watermarks to training or evaluation data via paraphrasing, similarly as what we study in this work. They exploit watermark radioactivity (Sander et al., 2024), i.e., the detectable traces left when watermarked text is used for training. Applications include detection of texts used in retrieval augmented generation (RAG) (Jovanović et al., 2025), benchmark contamination detection (Sander et al., 2025), and training data copyright (Zhang et al., 2025). Only Waterfall (Lau et al., 2024) evaluates post-hoc watermarking through LLM paraphrasing, focusing on code (MBPP) and natural text (C4 and arXiv) for provenance detection.
These works demonstrate the utility of such method for data protection but do not evaluate it as a general watermarking one and do not characterize failure modes across text types and settings.
3 Method
Post-hoc watermarking via paraphrasing.
The key idea is to paraphrase the input text using an LLM while applying watermarking during generation, as depicted in Figure 1. The pipeline is as follows:
  1. 1.
    Split input text into chunks (sentences or paragraphs).
  2. 2.
    For each chunk, prompt an LLM to produce a paraphrase, given specific instructions (e.g., preserve named entities, minimal lexical change) and some previous context.
  3. 3.
    During decoding, sample the next token with watermarking: for instance, favor tokens in a “greenlist” according to a strength parameter or any of the methods described in subsection 2.2.
  4. 4.
    Aggregate watermark scores across chunks and compute the detector statistic.
3.1 LLM Watermarking
Next token generation.
At each decoding step, the LLM computes logits $\bm{\ell}\in\mathbb{R}^{|\mathcal{V}|}$ over the vocabulary $\mathcal{V}$ conditioned on the preceding context. These logits are converted into a probability distribution $\mathbf{p}=\text{softmax}(\bm{\ell}/T)$ via temperature scaling by $T$.
Watermarked sampling.
Watermarking modifies the decoding process rather than sampling directly from $\mathbf{p}$. This is done depending on a window $\mathbf{w}$ of the $k$ preceding tokens and a secret key $s$. A pseudorandom function $\textsc{PRF}(\cdot)$ combines $\mathbf{w}$, $s$, and a candidate token $v$ to produce a value used for biasing the sampling of token $v$. For instance, in the green-red method (Kirchenbauer et al., 2023a), we compute $\textsc{PRF}(\mathbf{w},s,v)$ for all tokens $v\in\mathcal{V}$ to partition the vocabulary into a “greenlist” and a “redlist,” then bias the logits by adding $\delta$ to green tokens. Other schemes alter the sampling mechanism directly (e.g., with the Gumbel-max trick (Aaronson and Kirchner, 2023)). This embeds a statistical signal into the generated text that is imperceptible to readers but detectable by the detector. We provide a comprehensive description of the watermarking schemes evaluated in this work in subsection 6.2.
Detection and statistical test.
Detection recomputes the pseudorandom function for each group of tokens and aggregates the results into a test statistic. Under the null hypothesis $\mathcal{H}_{0}$ “not watermarked with this specific scheme”, the statistic should follow a known distribution, yielding a $p$-value for the probability of observing such a score or a bigger score by chance. Under the alternative hypothesis $\mathcal{H}_{1}$ “watermarked text with this specific scheme”, the statistic should deviate significantly enough and lead to low $p$-values. We flag text as watermarked if the $p$-value is below a detection threshold $\alpha$, which controls the false positive rate (FPR).
For instance, if $K$ (the test statistic here) is the observed number of green tokens out of $N$ scored tokens and that this statistic follows under $\mathcal{H}_{0}$ a Binomial distribution with parameters $N$ and $\gamma$ (expected green ratio, typically $0.5$), then $p\text{-value}=\mathbb{P}\left(X\geq K\mid X\sim\text{Binomial}(N,\gamma)\right)$. As an example, if we observe $K=65$ green tokens out of $N=100$ with $\gamma=0.5$, the $p$-value is approximately $10^{-3}$; to flag this text as watermarked, we would need a FPR of $\alpha\geq 10^{-3}$.
An important practical consideration is deduplication: because the pseudorandom function depends on the preceding $k$ tokens, repeated n-grams generate identical hashes. This violates the statistical independence of the scores used when assuming the distribution of the test statistic under $\mathcal{H}_{0}$. Aggregating scores only over unique watermark windows within the text (Kirchenbauer et al., 2023a; Fernandez et al., 2023) is a good way to mitigate this issue and to ensure valid $p$-values. But deduplication alone may not guarantee valid statistical tests because natural language inherently favors certain n-grams over others. This bias can cause $p$-values to be artificially low or high even for unwatermarked text. To address this, we ensure that under $\mathcal{H}_{0}$, $p$-values are approximately uniform on $U(0,1)$. In practice, we test many candidate secret keys and select one for which the empirical $p$-value distribution on unwatermarked text is close to uniform. We provide more details in subsection 6.1.
3.2 Compute-Driven Flexibility
Operating in a post-hoc setting offers more flexibility than in-generation watermarks, which must comply with live-serving constraints such as latency, memory, and decoding speed (Dathathri et al., 2024). Here, we can use large or small models, run higher temperatures or beam search instead of simple sampling, and even employ multi-candidate selection like WaterMax (Giboulot and Furon, 2024), as a function of available compute and quality targets. We detail the specific higher-compute decoding (and detecting) methods we test below.
Beam search for watermarking.
We use a beam search approach to improve the quality-detectability trade-off. Specifically, we maintain $B$ candidate sequences and expands each beam with $V$ candidates sampled from the watermarked probability distribution $\mathbf{p}_{\text{wm}}(\cdot\mid x_{<t})$, which is derived by applying the watermarking scheme (e.g., Maryland, SynthID, DIPMark, MorphMark) to the base model’s logits. To select the top-$B$ beams at each step, we score candidates using log-probability under a reference distribution. We explore two scoring variants: unbiased scoring uses the original model probabilities $\mathbf{p}_{\text{orig}}(\cdot\mid x_{<t})$ to favor sequences with lower perplexity relative to the base model, prioritizing text quality; biased scoring uses the watermarked probabilities $\mathbf{p}_{\text{wm}}(\cdot\mid x_{<t})$ to favor sequences that are most probable under the watermarked distribution, potentially yielding stronger watermark signals. We note that the latter was also explored for generation-time watermarking in (Kirchenbauer et al., 2023b). For candidate generation, we consider both deterministic beam search (selecting top-$V$ tokens by $\mathbf{p}_{\text{wm}}$) and stochastic beam search (sampling $V$ tokens from $\mathbf{p}_{\text{wm}}$).
Multi-candidate selection with Wateremax (Giboulot and Furon, 2024)
Rephrasing is done chunk by chunk of $L$ tokens. For each chunk, we generate $m$ candidates $\{\tilde{y}^{(1)},\dots,\tilde{y}^{(m)}\}$ without applying any logit bias. We then select the candidate that naturally maximizes the watermark score: $y^{*}=\arg\max_{y}\text{Score}_{\text{wm}}(y)$. This method is “distortion-free” regarding the sampling distribution but is typically infeasible in standard API usage due to cost (generating $L$ times more tokens per chunk).
Entropy-aware detection.
At detection time, we compute the entropy $H_{t}$ of the model’s predicted distribution at each token position $t$ in the watermarked text $H_{t}=-\sum_{v\in\mathcal{V}}p_{t}(v)\log p_{t}(v)$ where $p_{t}(v)$ is the model probability for token $v$ given the prefix up to t1t-1 in the watermarked text. We then apply an entropy filter: a token is included in the watermark score only if $H_{t}$ exceeds a chosen threshold $\tau$, similar to what is done in Lee et al. (2023) for code. This focuses detection on high-entropy positions, where watermarking is more effective, and ignores low-entropy tokens. All entropy computations are performed on the watermarked text, as the original text is not available at detection time. As a result, these entropy values differ from those at generation-time, when the model was conditioned.
4 Experiments
We conduct a comprehensive evaluation of post-hoc watermarking to assess: (1) detection power (the ability to reliably identify watermarked text, measured through $p$-values) and (2) fidelity (the preservation of meaning and quality through paraphrasing). In subsection 4.1, we detail the experimental setup, including datasets, models, watermarking schemes, decoding strategies, and evaluation metrics. In subsection 4.2, we compare watermarking methods on the quality-detectability Pareto frontier, finding that Gumbel-max dominates under standard sampling. In subsection 4.3, we show that larger models preserve semantics better, but small/mid-size models offer sweet spots, with stronger watermark signals due to higher entropy. In subsection 4.4, we demonstrate that beam search improves the Pareto frontier for all applicable methods. In subsection 4.5, we evaluate entropy-aware detection and find only modest gains. In subsection 4.6, we investigate post-hoc watermarking on code, revealing that correctness constraints limit detectability. In subsection 4.7, we evaluate cross-lingual robustness and, in subsection 4.8, the impact of chunking on long documents.
4.1 Experimental Setup
Datasets.
We evaluate on three diverse corpora to capture different text characteristics and use cases:
  • Books: Passages in English from the Gutenberg dataset (Project Gutenberg, 2025), containing books. These represent longer-form, literary text with complex sentence structures.
  • Wikipedia: Lead paragraphs from 500 randomly sampled Wikipedia articles. These represent factual, encyclopedic text with a formal writing style and different languages.
  • Code: Programming problems from the HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) benchmarks, consisting of Python functions with accompanying unit tests.
Paraphrasing models and prompts.
We experiment with several dense open-weight instruct LMs as paraphrasers, spanning different families and sizes to study the impact of model scale on paraphrase quality and watermark preservation. Specifically, we use the instruct version of LLaMA-3 (Dubey et al., 2024) (1B, 3B, 8B, 70B), Gemma-3 (DeepMind, 2025) (1B, 4B, 12B, 27B) , Qwen-2.5 (Team, 2024) (0.5B, 1.5B, 3B, 7B, 14B, 32B), and SmolLM-2 (Allal et al., 2025) (135M, 360M, 1.7B).
To ensure the models outputs the watermarked text directly and without any comments, we use a structured system prompt and prefill the assistant’s response with a prefix (e.g., “Rephrased text:”). An example prompt used is:
“You are a text rephrasing assistant. You must rephrase the given text while strictly preserving its original meaning, style, and structure. You must output only the rephrased text, with no explanations or commentary.”
Chunking strategy.
Large documents may be difficult to rephrase in a single pass, or may even exceed the model’s context window, so we split texts into chunks. We compare two modes: full-context, which processes the entire document at once, and context-aware chunking, which processes each chunk separately while prepending one or two previously rephrased chunks as context to preserve coherence.
Watermarking schemes.
We evaluate multiple watermarking methods (detailed in subsection 6.2), all with window size $2$, varying their hyperparameters to explore their detection-quality trade-offs.
  • Green-Red List (Kirchenbauer et al., 2023b): We vary the bias $\delta\in\{1.0,2.0,4.0\}$.
  • DiPMark (Wu et al., 2023): We vary the reweighting parameter $\alpha\in\{0.2,0.3,0.4\}$.
  • MorphMark (Wang et al., 2025): We fix $\kappa=10$ and vary $p_{0}\in\{0,0.05,0.1,0.2\}$.
  • SynthID-Text (Dathathri et al., 2024): We vary the number of tournaments $k\in\{10,20,30\}$. We also compute the $p$-values from weighted and non weighted scores as described in subsection 6.2.
  • Gumbel-max (Aaronson and Kirchner, 2023): Fixed by temperature and top-$p$.
Choice of private key.
We address the private key issue described in section 3 by testing 50 candidate keys. For each combination of tokenizer, watermarking method (and depth for SynthID), we evaluate the keys on non-watermarked texts to estimate the $p$-value distribution under $H_{0}$. We then select the key that minimizes the deviation from $U(0,1)$, as measured by the Kolmogorov–Smirnov statistic. Next, we verify that the theoretical FPRs for these keys match the empirical FPRs on 1.5M Wikipedia documents containing between 1k and 100k tokens (using watermark window deduplication). This ensures that the subsequent results under $H_{1}$ reflect the intrinsic performance of each scheme rather than artifacts of a particular key choice. Further details are provided in Table 6.
Entropy aware detection.
For all methods, we evaluate entropy-aware detection by computing per-token entropy with the same model used for rephrasing. At detection, we only score tokens with entropy $\geq\tau$, with $\tau\in\{0,0.2,0.4,0.6,\dots,1.8,2.0\}$. We note that we only pass the watermarked text to compute the per-token entropy without the original text as context.
Decoding strategies.
We explore different decoding approaches:
  • Nucleus sampling (Holtzman et al., 2019). We vary temperature $T\in\{0.7,1.0,1.2\}$ and top-$p\in\{0.9,0.95,0.99\}$. Higher temperature or top-$p$ should increase entropy at the cost of faithfulness.
  • Beam search (Sutskever et al., 2014), as detailed in section 3. We test beam widths $B\in\{3,5,10\}$ with same number of candidates per beam. We experiment with selection criteria based on either the non-watermarked likelihood or the watermarked model’s likelihood, and use stochastic (temperature $1$ and top-$p$ 0.95) or deterministic selection within beams.
  • WaterMax (Giboulot and Furon, 2024), as detailed in section 3. We generate $L\in\{4,8,16\}$ non-watermarked candidate tokens per draft $m\in\{4,8\}$ and select the one with the highest number of green tokens, and repeat.
Evaluation metrics.
We employ the following metrics to evaluate performance:
  • Detection: We compute the $p$-value of the corresponding detection test, which can be read as the probability that a watermark score at least as high would happen if the text was not watermarked.
  • Quality: We evaluate rephrase quality along three axes. Semantic similarity: We use the BERTScore (Zhang et al., 2020) to measure meaning preservation between original and rephrased texts. Naturalness: We compute the average cross entropy of the rephrased text using Mistral-7B-v0.3 (Jiang et al., 2023), conditioned on the prefix “This is a rephrased version of [ORIGINAL TEXT]:”, capturing if the output is fluent and plausible. Length ratio: We report the character-level length ratio between rephrased and original text, which should remain close to $1$.
  • Functional Correctness (Code): For code datasets, we measure the percentage of paraphrased functions that pass their unit tests (Pass@1).
In Sections 4.2, 4.3, 4.5, 4.4 and 4.8, each reported value corresponds to the median over 100 passages of approximately 800 characters ($\sim$200 tokens) sampled from Charles Dickens’ novels in the Project Gutenberg dataset (Project Gutenberg, 2025). We strictly retain only rephrasings where the output length remains comparable to the original (length ratio $\in[0.75,1.25]$), reporting results only for settings where at least 70% of passages satisfy this constraint.
As detailed in section 8, output length distributions are remarkably consistent across watermarking schemes and are primarily determined by the base model rather than the watermarking method. While output length is arguably a quality signal in itself, we invite the reader to examine these distributions in the appendix, allowing the main analysis to focus on semantic fidelity and detection power within valid generation bounds.
Table 1: Qualitative comparison on literary text. A sample from A Christmas Carol by Charles Dickens, rephrased using Llama-3.2-3B-Instruct with the Gumbel-max scheme, $T=0.7$ and top-p = 0.95. The output maintains the original tone and structure while achieving high semantic similarity and good watermark detectability.
Original Text Watermarked Rephrasing
“You don’t believe in me,” observed the Ghost. “I don’t,” said Scrooge. “What evidence would you have of my reality beyond that of your senses?” “I don’t know,” said Scrooge. “Why do you doubt your senses?” “Because,” said Scrooge, “a little thing affects them. A slight disorder of the stomach makes them cheats. You may be an undigested bit of beef, a blot of mustard, a crumb of cheese, a fragment of an underdone potato. There’s more of gravy than of grave about you, whatever you are!” Scrooge was not much in the habit of cracking jokes, nor did he feel, in his heart, by any means waggish then. The truth is, that he tried to be smart, as a means of distracting his own attention, and keeping down his terror; for the spectre’s voice disturbed the very marrow in his bones. To sit, staring at those fixed glazed eyes, in silence for a moment, would play, Scrooge felt, the very deuce with him. “You’re not convinced of my existence,” said the Ghost. “I am not,” replied Ebenezer Scrooge. “What evidence do you have to confirm my reality beyond the limits of your perceptions?” “I don’t know,” Scrooge replied. “Why do you question the reliability of your senses?” “Because,” Scrooge said, “a minor aberration can significantly impact them. A slight irregularity in digestion can cause them to be misled. You may be an undigested morsel of meat, a splash of mustard, a fragment of a partially cooked potato. There’s more of gristle than of flesh about you, regardless of what you are!” Scrooge was not accustomed to indulging in humor, nor did he feel that he was particularly witty at any point. The truth was, that he tried to appear intelligent, as a means of diverting his own attention and suppressing his fear; as a way to distract himself from the unsettling presence of the spectre. To sit, staring at those fixed, unblinking eyes, in silence for a moment, would play, Scrooge felt, a cruel and unforgiving reminder of the truth.
Stats: SBERT: 0.904   $\bullet$  Ratio: 1.07   $\bullet$  Perplexity: 2.61   $\bullet$  $p$-val: $1.7\times 10^{-5}$
4.2 Quality-Detection Trade-Off
We first evaluate the fundamental trade-off between rephrasing quality and detection strength in Figure 2, using Llama-3.2-3B-Instruct with random sampling. Each point corresponds to a distinct parameter configuration as described in subsection 4.1. Optimal methods occupy the top-right region of the plot, combining high text quality with strong detectability. The Gumbel-max method appears to dominate this Pareto frontier, while DiPMark and MorphMark lag behind.
Figure 2: Pareto fronts of watermarking methods showing the trade-off between quality and watermark strength using Llama-3.2-3B-Instruct. Each point corresponds to a different parameter configuration, with values representing medians across 100 rephrased passages. Experimental details are given in Sections 4.1 and 4.2.
We also present a qualitative example in Table 1, where we rephrase text using Llama-3.2-3B-Instruct with the Gumbel-max watermark with $T=0.7$ and $p=0.95$. The rephrasing maintains the original tone, structure, and factual content. Additional examples are provided in Table 7 of section 7.
4.3 Impact of Model Family and Scale
We investigate whether larger models are better suited for post-hoc watermarking. Figure 3 shows the trade-off curves by model family (SmolLM, Gemma, Llama, Qwen) and size. Regarding performance, we observe that larger models preserve semantics better: across all families, increasing model size (darker points) shifts the clusters upward. For example, within the Llama family, Llama-3.1-8B (dark green) attains lower cross entropy than the 3B and 1B variants (lighter teal) at comparable watermark strengths. However, larger models struggle to reach low $p$-values compared to smaller models: they populate the left rather than the right of the plots. Gemma 3 models are essentially absent from the high-detectability region, as their outputs naturally exhibit very low entropy, even for the smallest sizes. Overall, when strong watermarking is required, only small models appear on the frontier, whereas larger models dominate in the lower-strength regime. We see in subsection 4.6 that the issue with large models persists even if temperature is increased.
Note that these results are conditioned on the length filtering described in subsection 4.1; thus, the presence of a data point implies enough successful valid generation for the corresponding combination of model/scheme/parameters. We refer the reader to section 8 to assess the generation stability of each configuration, as smaller models fail to meet length constraints more frequently, cannot be read from the figure.
Figure 3: Impact of model family and size. Cross Entropy vs. watermark strength. Larger models improve quality for a given watermark strength, but small models are necessary to reach high strengths. All families are comparable, except Gemma-3 that is not suitable. Experimental details are given in Sections 4.1 and 4.3.
4.4 Decoding Strategies: Beam Search vs. Sampling vs. WaterMax
We explore whether systematic search can find better watermarked sequences than random sampling. Figure 4 compares standard sampling (crosses) with beam search for suitable methods: Green-red, Synth-ID, MorphMark, and DiPMark.
We observe that beam search consistently improves the Pareto frontier, especially with biased scoring (see section 3): it shifts the results upward, substantially improving rephrasing quality at a fixed watermark strength. Notably, while Gumbel-max sampling appears to be the best option when used with random sampling in Figure 2, other methods can be substantially improved through beam search.
WaterMax (Giboulot and Furon, 2024) results are shown in maroon in Figure 2. Surprisingly, we found that WaterMax achieved weak detectability across the hyperparameters we tested, making it difficult to achieve strong watermarking guarantees.
Figure 4: Beam search improves the Pareto frontier. Cross entropy vs. watermark strength for suitable methods. Beam search, especially with biased scoring (see section 3), shifts the frontier upward, substantially improving rephrasing quality at a fixed watermark strength. Experimental details are given in Sections 4.1 and $4.$
4.5 Entropy-Aware Detection
Instead of scoring every token, we filter out low-entropy tokens at detection time, as motivated in section 3 and detailed in subsection 4.1. In Figure 5, we fix the rephrasing model to Llama-3.2-3B and, for each method and watermarking configuration, test whether there exists an entropy threshold that improves detectability by more than $5\%$ on at least 50 of the 100 texts of Charles Dickens.
The left plot reports, for each watermarking method, the proportion of configurations for which such a threshold exists. We see that most methods have only 30304040 successful configurations. The middle plot shows, for those successful configurations, the actual gain in detection performance, which never exceeds $20\%$. The right panel presents the corresponding non-aggregated statistics.
We see that WaterMax is different: any entropy threshold degrades detectability. This is because WaterMax optimizes the watermark score at the sentence level rather than the token level; filtering by token entropy does not expose additional signal, it only reduces the number of scored tokens and thus the available statistical evidence. One other and related property of WaterMax is that it cannot be used for active dataset inference through radioactivity. This is explained in details in section 8.
Figure 5: Effect of entropy-aware detection. Left: Share of configurations for which some threshold improves detection by more than $5\%$ on at least half of the texts. Middle: Median relative improvement for those configurations. Right: For every configuration, fraction of texts helped at optimal threshold vs. median improvement; dashed lines mark the $50\%$ and $5\%$ criteria. Experimental details are given in Sections 4.1 and 4.5.
4.6 Post-hoc Code Watermarking
Experimental details.
We now conduct experiments on HumanEval and MBPP to evaluate the feasibility of post-hoc watermarking on structured and verifiable text, where watermarking requires to preserve the syntax and the function of the code.
Data preparation and preprocessing. For each problem instance, we concatenate the provided prompt (e.g., function signature and docstring) and the canonical solution. We make sure that the last top-level function definition of the code is the target for execution by moving it if necessary. We show in App. 7.1 an example of a HumanEval task. Overall, this dataset comprises 164 HumanEval and 974 MBPP code samples, with average lengths of 180 and 80 tokens, respectively. The code samples range from approximately 20 to 500 tokens in length.
Watermarking pipeline. The watermarking process is applied only to the code (we exclude unit tests to preserve the evaluation). Since watermarking may inadvertently rename function identifiers, we implement a post-processing step that parses the output and restores the original entry point name required by the benchmark test harness. At test time, we concatenate the watermarked code with the test harness to evaluate functional correctness.
Metrics for code watermarking.
First, we measure functional correctness: the fraction of watermarked code that passes the tests (pass@1). If pass@1 is low, one could regenerate with different random seeds until obtaining functional code, but this assumes tests are available and it is cumbersome; ideally, pass@1 should be high. Second, as before we measure the detection power: the true positive rate (TPR) at a fixed false positive rate (FPR), and/or the distribution of $\log_{10}$ $p$-values. These can be computed among all rephrased code, or more interestingly among only the ones that passed the tests. In practice, the latter is a better indicator of the watermark power, since codes that do not pass tests often exhibit degenerative patterns (repeated code or text at the end) which artificially increase the TPR.
Table 2: Qualitative example of code watermarking on a sample from MBPP (using Llama-3.1-8B at temperature 1.4 with Gumbel-max watermarking, top-$p=0.95$). The post-hoc rephrasing here refactors variable names to ensure detection (low p-value) while maintaining correctness.
Original Text Watermarked Rephrasing
# A python function to identify non-prime numbers.
import math
def is_not_prime(n):
 result = False
 for i in range(2,int(math.sqrt(n)) + 1):
   if n % i == 0:
    result = True
 return result
# A python function to identify non-prime numbers
import math
def is_not_prime(number):
 verdict = False
 for divisor in range(2, int(math.sqrt(number)) + 1):
   if number % divisor == 0:
    verdict = True
 return verdict
$p_{\text{value}}$: $0.12$   $\bullet$  Tokens: $48$   $\bullet$  Correct: ✓ $p_{\text{value}}$: $3.29\times 10^{-3}$   $\bullet$  Tokens: $57$   $\bullet$  Correct: ✓
Comparing watermarking methods.
[…]
Comparing models.
Figure 7: Evaluation of post-hoc watermarking accross models using Gumbel-max (top-pp=0.95). (a) Distribution of -$\log_{10}$ p-value for watermarked codes that pass the functional correctness test (bars indicate mean and median). Different colors represent model families, with darker shades indicating larger models. Corresponding values for Llama models are detailed in Table 3. (b) Pass@1 and detection power (median -$\log_{10}$ p-value among correct codes) for Llama 3B, 8B and 70B models across temperatures. Different model families and sizes behave differently in terms of detection power at fixed temperatures, and interestingly, smaller models can achieve better Pareto optimality than larger models.
We compare model families and sizes using Gumbel-max watermarking (top-pp=0.95). Figure 7(a) shows that larger models lead to lower detection power (lower -$\log_{10}$ p-values), as their outputs are less entropic and leave less room for watermarking. Table 3 confirms this: Llama-70B achieves median -$\log_{10}$ p-values around 0.5-0.7, while smaller models reach values near 2 at $T=1.4$. However, higher detection power in smaller models comes at the cost of functional correctness; for example, at $T=1.4$, Llama-8B achieves a TPR of 0.29 among correct codes but a pass@1 of only 0.71.
To further investigate this trade-off, we sweep the temperature in Figure 7(b). Interestingly, the smaller 8B model achieves a better Pareto frontier than the larger 70B model, attaining higher detection power at equivalent pass@1 levels. This suggests that while larger models produce higher-quality code, their reduced entropy limits watermark effectiveness, making smaller models more suitable for post-hoc code watermarking when balancing functionality and detectability.
[…]
$T=1.0$ $T=1.2$ $T=1.4$
[…]
All Passed All Passed All Passed All Passed All Passed All Passed
Llama-3.2-1B-Instruct 0.59 0.60 0.51 0.03 0.01 0.44 1.32 0.88 0.22 0.07 0.16 9.10 1.56 0.72 0.29
Llama-3.2-3B-Instruct 0.81 0.60 0.56 0.01 0.01 0.73 0.87 0.75 0.08 0.05 0.53 1.84 1.00 0.34 0.13
Llama-3.1-8B-Instruct 0.92 0.74 0.74 0.04 0.03 0.89 1.16 1.11 0.12 0.11 0.71 2.31 1.83 0.41 0.29
Llama-3.3-70B-Instruct 0.92 0.38 0.37 0.00 0.00 0.92 0.41 0.40 0.00 0.00 0.90 0.44 0.44 0.00 0.00
4.7 Multi-Lingual Robustness on Wikipedia
To evaluate the generalization of post-hoc watermarking beyond English, we apply the method to paragraphs of Wikipedia articles in multiple languages. In the following, we use Llama-3.1-8B-Instruct, with Gumbel-max watermarking with varying temperature settings (0.8, 1.0, 1.2) and fixed top-$p=0.95$. Table 4 compares the performance across English (en), Spanish (es), French (fr), and Russian (ru). To not conflate the effect of number of tokens used for scoring with language, we only consider outputs that contain between 400 and 600 tokens, and aggregate results over 1k samples.
Table 4: Evaluation of post-hoc watermarking for different languages. We report the median log10 p-value and the SBERT score for semantic similarity, at varying temperature of the generation. While not impossible, watermarking languages other than English comes at a steeper cost on semantic quality.
Temperature en es fr ru
$\log_{10}p$ SBERT $\log_{10}p$ SBERT $\log_{10}p$ SBERT $\log_{10}p$ SBERT
0.80 1.832 0.960 1.608 0.949 0.810 0.946 1.355 0.908
1.00 3.813 0.954 4.408 0.940 2.809 0.937 4.976 0.888
1.20 8.147 0.946 20.098 0.921 19.705 0.911 23.689 0.836
The results indicate a performance gap: while non-English languages (Spanish, French, Russian) can be watermarked, they suffer from a steeper trade-off. Achieving high detection strength in these languages requires a larger sacrifice in semantic quality compared to English. This suggests that the logits of the paraphrasing model are less easy to manipulate in languages for which the model was not primarily trained, since most training data is in English. It may also reflect the fact that our semantic similarity metric (SBERT) is more accurate for English than for other languages.
4.8 Impact of Chunking Strategy on Long Documents
To assess the impact of chunking as described in section 3), we compare full-context processing with context-aware chunking (500-token chunks, with up to 1000 context tokens from previously rephrased chunks) on documents of varying length (500–4000 tokens), using Dickens novel excerpts and Llama-3.2-3B-Instruct.
For each document length, results are averaged over 5 independent texts. Table 5 shows that full-context processing increasingly summarizes the content, while chunking better preserves length and yields stronger detection at better semantic similarity. This indicates that context-aware chunking is crucial for reliable watermarking of important document lengths.
Size Length Ratio Detection ($\log_{10}p$) Similarity
Full Chunked Full Chunked Full Chunked
500 0.91 0.91 19.3 19.3 0.94 0.94
1500 1.15 0.97 49.5 148.8 0.85 0.91
2500 0.78 0.98 56.8 195.5 0.91 0.91
4000 0.70 0.86 24.9 198.3 0.84 0.91
Table 5: Context-aware chunking vs. full-context. Comparison of full-context and context-aware chunked rephrasing (500-token chunks with up to 1000 context tokens) on documents of varying length, averaged over 5 excerpts per length. Experimental details are given in subsection 4.1 and subsection 4.8.
5 Conclusion
We presented a comprehensive evaluation of post-hoc text watermarking via LLM rephrasing, a paradigm that enables embedding traceable signals into existing text. Unlike generation-time constraints, the post-hoc setting allows the allocation of additional compute to optimize the trade-off between text quality and watermark detectability.
Our experiments yielded several findings. First, the simplest Gumbel-max scheme (Aaronson and Kirchner, 2023) achieved better trade-offs than all other tested methods under random sampling. We note that in the generation-time watermarking literature, Gumbel-max is criticized for its deterministic nature: it fixes the randomness of generation, so identical prompts produce identical outputs, which can be problematic in production. In our post-hoc setting, however, this limitation is less consequential. We also observe that all other schemes benefit substantially from beam search. Second, to achieve high watermark strength, smaller, more entropic models outperformed larger models run at high temperature. Third, entropy filtering at detection time provided only marginal gains while introducing additional complexity, making it a less practical option in typical deployment scenarios.
A limitation of our study, and the field at large, is the reliance on automated metrics such as perplexity and BERTScore to approximate text quality and semantic preservation. Much like in image and audio watermarking, these proxies cannot fully capture the nuances of human perception, particularly across different languages where robust human evaluations remain essential. To address this limitation, we extended our analysis to post-hoc code watermarking, where execution-based correctness provides an objective, ground-truth measure of utility. Our results reveal that while watermarking is feasible in this domain, and that the Gumbel-max scheme also prevails, strict correctness constraints reduce the available capacity for watermark embedding compared to natural language.